# RAPiD-Seg: Range-Aware Pointwise Distance Distribution Networks for 3D LiDAR Segmentation

### v1

@@ 1 Introduction
As an enabler, we propose the Range-Aware Pointwise Distance Distribution (RAPiD-Seg) solution for LiDAR segmentation. It utilizes invariance of RAPiD features to rigid transformations and point cloud sparsity variations, while concentrating on compact features within specific local neighborhoods. Specifically, our method leverages inherent LiDAR isotropic radiation and semantic categorization for enhanced local representations while reducing computational burdens. Moreover, our formulation computes a 4D distance, incorporating both 3D geometric and reflectivity differences to enhance semantic segmentation fidelity. To compress high-dimensional RAPiD features into tractable voxel-wise embeddings, we propose a novel embedding approach with our class-aware double-nested AutoEncoder (AE) module. We further incorporate a channel-wise attention fusion and a two-stage training strategy to optimize the AE independently before integrating into the full network, enhancing performance and generalization capability.
@@ 4 RAPiD Embedding
$\mathcal{L}_{\text{contr}}$ $=\frac{1}{m}\sum_{i=1}^{m}\left[{\footnotesize\frac{1}{|P(i)|}}\sum_{p\in P(i)}{\footnotesize\text{ReLU}(\alpha-\text{sim}(H_{i},H_{p}))}\right.$
$\quad\quad+\left.{\footnotesize\frac{1}{|N(i)|}}\sum_{n\in N(i)}{\footnotesize\text{ReLU}(\text{sim}(H_{i},H_{n})-\alpha)}\right],\vspace{-8pt}$ (15)
@@
where $i$ represents the point index; $P(i)$ and $N(i)$ are the indices of the nearest point in the same (Positive) and different class (Negative) relative to point $i$, respectively; $\text{sim}(\cdot)$ is a similarity measure; $\alpha$ is a margin controlling separation between classes. The pointwise representations $H_{i}$, $H_{p}$, and $H_{n}$ are broadcasted from $\hbar$ based on voxel indices $I^{v}$.
@@ 6.3 Ablation Studies
Table 4: Effects of RAPiD and Reflectivity features compared to other configurations on SemanticKITTI validation set. PDD: the compressed PDD embeddings; RAPiD-R: only RAPiD, without reflectivity feature fusion; RAPiD-R: both RAPiD and reflectivity.
@@ 7 Conclusion and Discussion
Our RAPiD features hold significant potential for application in various tasks such as point cloud registration, semi-/weakly- supervised learning, and extend to promising applications in multi-modal research in the future.

### v3

@@ 1 Introduction
As an enabler, we propose the Range-Aware Pointwise Distance Distribution (RAPiD-Seg) solution for LiDAR segmentation. As shown in Fig. 1, it utilizes invariance of RAPiD features to rigid transformations and point cloud sparsity variations, while concentrating on compact features within specific local neighborhoods. Specifically, our method leverages inherent LiDAR isotropic radiation and semantic categorization for enhanced local representations while reducing computational burdens. Moreover, our formulation computes a 4D distance, incorporating both 3D geometric and reflectivity differences to enhance semantic segmentation fidelity. To compress high-dimensional RAPiD features into tractable voxel-wise embeddings, we propose a novel embedding approach with our class-aware double-nested AutoEncoder (AE) module. We further incorporate a channel-wise attention fusion and a two-stage training strategy to optimize the AE independently before integrating into the full network, enhancing performance and generalization capability.
@@ 4 RAPiD Embedding
$\mathcal{L}_{\text{contr}}$ $=\frac{1}{m}\sum_{i=1}^{m}\left[{\frac{1}{|P(i)|}}\sum_{p\in P(i)}{\text{ReLU}(\alpha_{p}-\text{sim}(H_{i},H_{p}))}\right.$
$\quad\quad+\left.{\frac{1}{|N(i)|}}\sum_{n\in N(i)}{\text{ReLU}(\text{sim}(H_{i},H_{n})-\alpha_{n})}\right],$ (15)
@@
where $i$ represents the point index; $P(i)$ and $N(i)$ are the indices of the nearest point in the same (Positive) and different class (Negative) relative to point $i$, respectively; $\text{sim}(\cdot)$ is a similarity measure; $\alpha_{p}$ and $\alpha_{n}$ is the positive and negative margin controlling separation between classes, respectively. The pointwise representations $H_{i}$, $H_{p}$, and $H_{n}$ are broadcasted from $\hbar$ based on voxel indices $I^{v}$.
@@ 6.3 Ablation Studies
Table 4: Effects of RAPiD and Reflectivity features compared to other configurations on SemanticKITTI validation set. PDD: the compressed PDD embeddings; RAPiD-R: only RAPiD, without reflectivity feature fusion; RAPiD+R: both RAPiD and reflectivity.
@@ 7 Conclusion and Discussion
Our RAPiD features hold significant potential for application in various tasks such as object detection [30], point cloud registration, semi-/weakly- supervised learning, and extend to promising applications in multi-modal research in the future.